# references.bib

### v3

@@ 6 Conclusion
We then established that the power-law tailed distribution of metaorder durations is not sufficient to counteract impact decay, leading to price sub-diffusion. Rather, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders, which is indeed present in data. In fact, we conjecture that the intra- and cross-correlations between child orders decay roughly in the same manner, a feature that may be crucial for explaining the success of the construction of synthetic metaorders from public data [maitrier2025generatingrealisticmetaorderspublic].
The existence of correlations between metaorders is therefore a crucial ingredient to recover price diffusion. The old debate between order splitting and “herding” that seemed to have been closed by several papers in favor of splitting [toth2015equity, sato2023inferring], is perhaps not so clear-cut. Such correlations could be due either to the fact that many participants use the same trading signals, or that copy-cat metaorders follow past order flow, or else that some traders successfully predict the future behaviour of other participants. Note that within our story any predictive alpha signal manifests itself through autocorrelated metaorders, since only trades can move the price (see also the discussion in [bouchaud2018trades], chapter 20).
@@
Such noteworthy agreement between theory and data suggests that our framework correctly captures the basic mechanism at the heart of price formation, namely the average impact of metaorders. We claim that our results strongly support the “Order-Driven” theory of markets, according to which it is the mechanical impact of trades, independently of any notion of “fundamental information”, that generates volatility in financial markets, a picture advocated in [bouchaud2006random, hopman2007supply, bouchaud2018trades, bouchaud2022inelastic] and, in a different context, in [gabaix2021search]. In particular, the Efficient Market Hypothesis, which posits that volatility is mostly due to the variation of “fundamental value”, cannot easily explain our results on the correlations between price changes and order flow imbalance.
In a follow-up paper (G. Maitrier et al., in preparation), we will confirm these ideas using numerical simulations of our model, in order to (i) confirm the qualitative validity of our theoretical analysis, ii) propose a efficient way to generate an exhaustive dataset, capturing the subtle interplay between order flows and prices, and (iii) validate on realistic synthetic data the method of metaorder reconstruction from the public tape recently proposed in [maitrier2025generatingrealisticmetaorderspublic].

### v4

@@ 6 Conclusion
We then established that the power-law tailed distribution of metaorder durations is not sufficient to counteract impact decay that leads to price sub-diffusion. Rather, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders, which is indeed present in data. In fact, we conjecture that the intra- and cross-correlations between child orders decay roughly in the same manner, a feature that may be crucial for explaining the success of the construction of synthetic metaorders from public data [maitrier2025generatingrealisticmetaorderspublic].
The existence of correlations between metaorders is therefore a crucial ingredient to recover price diffusion. The old debate between order splitting and “herding” that seemed to have been closed by several papers in favor of splitting [toth2015equity, sato2023inferring], is perhaps not so clear-cut. Such correlations could be due either to the fact that many participants use the same trading signals, or that copy-cat metaorders/unwinding market maker inventories correlate to past order flow, or else that some traders successfully predict the future behaviour of other participants. Note that within our story any predictive alpha signal manifests itself through autocorrelated metaorders, since only trades can move the price (see also the discussion in [bouchaud2018trades], chapter 20).
@@
Such noteworthy agreement between theory and data suggests that our framework correctly captures the basic mechanism at the heart of price formation, namely the average impact of metaorders. We claim that our results strongly support the “Order-Driven” theory of markets, according to which it is the mechanical impact of trades, independently of any notion of “fundamental information”, that generates volatility in financial markets, a picture advocated in [bouchaud2006random, hopman2007supply, bouchaud2018trades, bouchaud2022inelastic] and, in a different context, in [gabaix2021search]. In particular, the Efficient Market Hypothesis, which posits that volatility is mostly due to the variation of “fundamental value”, cannot easily explain our results concerning the correlations between price changes and order flow imbalance.
All these ideas will be quantitatively confirmed in a follow-up paper (G. Maitrier et al., in preparation) using a numerical simulation of our model. Such a numerical implementation will allow us in to (i) generate realistic synthetic datasets that capture the subtle interplay between order flows and prices, and (ii) validate on such synthetic data the method of metaorder reconstruction from the public tape recently proposed in [maitrier2025generatingrealisticmetaorderspublic].